OpenStudy (anonymous):

P = 2w + 2h what is w.

OpenStudy (anonymous):

Subtract 2h from both sides P-2h = 2w Divide both sides by 2 P-2h / 2 = w

OpenStudy (anonymous):

so w = P - 2h - 2 ?

OpenStudy (anonymous):

(P-2h)/2

OpenStudy (anonymous):



OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

can you do one more for me ?

OpenStudy (anonymous):

Welcome : )

OpenStudy (anonymous):

y varies indirectly as twice x. When x = 1, y = 3. Find y when x = 4.